ThF^{+} AS A CANDIDATE FOR eEDM MEASUREMENTS

\underline{MICHAEL C. HEAVEN}, JOSHUA H. BARTLETT, Department of Chemistry, Emory University, Atlanta, GA 30322.

The low-lying ^{3}Δ_{1} state of ThF^{+} has advantageous properties for studies of the electron electric dipole moment. Calculations indicate that internal fields as high as 90 GV/cm can be generated, and there is just one isotope of ThF with significant natural abundance. Previous experiments show two low-lying electronic states of ThF^{+} spaced by 315 cm^{−1} (JCP 136, 104305 (2012)). These are ^{1}Σ and ^{3}Δ_{1}, but the ordering is still in question. Experiments are in progress to determine the electronic angular momentum of the ground state. Laser excitation spectra for ThF^{+} have been recorded for transitions in the 19300-21100 cm^{−1} range. For eEDM studies it is technically more convenient to work with lower energy transitions. Survey spectra covering the range 15000-19000 cm^{−1} are currently being recorded. Results from these measurements will be presented and their relevance to possible eEDM studies will be discussed.